find the equation of straight line ehich passes through the points 3,5 and makes x intercept double then y intercept



1
Expert's answer
2021-07-05T16:22:11-0400

line crossing positive x axis and y axes. 2 points "(0, k) and (2k, 0)" or"-2" slope.

"Ax + By = C" has slope of "\frac{-A}{B}" .

Therefore, "2x + y = C."

"2*3 + 5 = 11."

Hence "2x + y = 11" .

x-intercept will be "\frac{11}{2}" and y-intercept will be "11" .

Therefore the equation will be

"y = -2x + 11"